How many grams of Al2O3 are formed from 15.0 L of O2 at 97.3 kPa & 21°C? 


Al + O2 = Al2O3

1
Expert's answer
2021-11-05T01:06:53-0400

Using PV=nRT

n=PV/RT

n=(97.3×15)/8.31×294

n=0.6

Mass =mole×molar mass

Mass =0.6×32

Mass=19.2g

If 32g of O2 produces 102g of Al2O3

Then 19.2g will produce Xg of Al

X=61.2g of Al2O3.